Hello, thanks for the reply. I've wasted too much time as it is trying to solve it...as this section of this book is VERY poor on examples...Look for my review on amazon.com when I'm finished with it...for now can someone please show me the way?
Question 1040677: hello, can you give me the solution set of ordered pairs to
y=5x-3 and y=x+5 and show how you arrived at the answer?
thank you much
p.s. their answer is (2,7)
Answer by ikleyn(5730)

.
hello, can you give me the solution set of ordered pairs to
y=5x-3 and y=x+5 and show how you arrived at the answer?
thank you much
p.s. their answer is (2,7)
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Since the two equations
y = 5x-3 and
y = x+5
have identical left sides, their right sides are equal too:
5x-3 = x+5.
I am sure you can easily solve it on your own and find x.
Then find y.
In this way, you will get the answer.

Found 2 solutions by JoelSchwartz, ikleyn:
Answer by JoelSchwartz(130)

y=5x-3
y=x+5
5x-3=x+5
5x-x-3+3=x-x+5+3
4x=8
4x/4=8/4
x=2
y=2+5
y=7

Answer by ikleyn(52874)

.
What?

You have an assignment to solve the system of two linear equations
and you can not solve simplest linear equation?
Single linear equation in one unknown?


Shocking ! Positively shocking !!